Lisa has scored
64
,
86
, and
84
on her previous three tests. What score does she need on her next test so that her average (mean) is
79
?

Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

(64+86+84+x)/4=79
234+x=79*4
x=316-234
x=82

Answer by Alan3354(69443)

Lisa has scored
64
,
86
, and
84
on her previous three tests. What score does she need on her next test so that her average (mean) is
79
?
===========
64 + 86 + 84 = 234
---
79*4 = 316
316 - 234 = what she needs.